# Ground-Compose-Reinforce: Tasking Reinforcement Learning Agents through Formal Language

### v1

@@ 3. RQ3
    Reinforcement Learning: Can agents use these grounded interpretations to autonomously learn new tasks through RL with self-generated rewards?
  4. RQ4
    Sparsity: Does our potential-based reward shaping strategy enable RL in sparse settings?
@@ 7.1 Experimental Setup
Figure 3: DrawerWorld (right) is a MuJoCo robotics environment where the agent can interact with two drawers and three boxes. Propositions capture whether each drawer is open, each box is lifted by the agent, and if a given box is in a given drawer.
@@
We evaluated our approach on two domains: an image-based gridworld with randomized object locations called GeoGrid (introduced in our running example), and a continuous-control MuJoCo robotics environment called DrawerWorld (Figure 3). Full environment details are provided in Appendix B.1.
Grounding Datasets. To emulate realistic pretraining conditions, we collected the grounding dataset $\mathcal{D}$ under minimal assumptions about downstream tasks. In GeoGrid, we generated trajectories using a random policy. In DrawerWorld, we manually controlled the robot to perform generic behaviours (e.g., opening drawers, picking up boxes) under the constraint that each trajectory interacted with at most one box. We labelled these trajectories using a hand-crafted labelling function over a fixed vocabulary $\mathcal{AP}$. Note that such a labelling function is not a hard requirement—labelling could be done through other means such as human annotation or contrastive learning.
Tasks. We designed a diverse set of RM tasks (Table 1). To evaluate RQ2, we included tasks that involve behaviours rarely or never seen in $\mathcal{D}$: e.g., in DrawerWorld, the Pickup-Each-Box task requires interacting with all three boxes, whereas no single training trajectory involves interaction with more than one.
@@
Table 1: List of RM tasks, with a description for each task. Tasks labelled (g) involve a temporally extended goal, tasks labelled (s) involve safety constraints, and tasks labelled (r) involve repeating some behaviour as often as possible. For each task, we report the mean and max (undiscounted) return across all trajectories in $\mathcal{D}$, along with the highest possible average return for any policy (or if unknown, we report the highest average return we are aware of, denoted with $\sim$). Some tasks involve behaviours that are rarely or never observed in $\mathcal{D}$.
Task Description Return in $\mathcal{D}$ Max Possible
Mean Max Return
@@
Sequence (g) Go to a red $\triangle$, then a green $\triangle$, then a blue $\triangle$. $0.04$ $1$ $1$
Loop (r) Repeatedly go to a red $\triangle$, then a green $\triangle$, then a blue $\triangle$. $0.04$ $3$ $\sim 5.36$
Logic (g+s) Go to all six objects, but always go to red objects before blue objects, and blue objects before green objects. $0.00$ $1$ $1$
Safety (g+s) Go to a red object, then a blue object, then a green object, but always avoid $\triangle$. $-0.84$ $1$ $1$
@@
Lift-Red-Box (r) Lift and hold the red box as long as possible. $41.7$ $736$ $\sim 1538$
Pickup-Each-Box (g) Pick up the red box, then the blue box, then the green box. 00 00 $1$
Show-Green-Box (g) Reveal the green box if it’s in a closed drawer, then lift it. $0.22$ $1$ $1$
7.2 Method and Baselines
Ground-Compose-Reinforce (GCR). We implemented the full version of Ground-Compose-Reinforce with potential-based reward shaping. Both the labelling function $\hat{\mathcal{L}}$ and PVFs are neural networks trained from the fixed grounding dataset $\mathcal{D}$. In GeoGrid, we trained PVFs using Fitted Q-Iteration [70]. In DrawerWorld, we trained PVFs to predict Monte Carlo returns. [^3] We also evaluated MCQ [71], a state-of-the-art offline RL method, but it performed worse than Monte Carlo regression. We attribute this to the relatively small size of $\mathcal{D}$ compared to standard offline RL benchmarks.
Baselines. To answer RQ1-RQ3, we compared against the following end-to-end (i.e. non-compositional) baselines. Since these baselines do not use RMs to model temporal dependencies, they instead condition on past observations using a GRU [72]. See Appendix B.2 for complete details.
LTL-conditioned Behaviour Cloning (LTL-BC) is a neural network policy that directly maps LTL formulas to behaviours. We labelled each trajectory $\tau$ in $\mathcal{D}$ with a description $\varphi$ in LTL (based on the propositional labels for $\tau$ in $\mathcal{D}$), then trained the policy to maximize the log-likelihood of actions in $\tau$, conditioned on $\varphi$. For each downstream task in Table 1, we evaluated the policy by conditioning on an LTL formula that approximately aligns with that task. We also considered bespoke models trained with advance knowledge of the downstream tasks in Table 1. Bespoke Reward Model directly predicts rewards, optimal values, and terminations for all downstream tasks simultaneously, enabling RL with shaped rewards. To train this model, we labelled each trajectory $\tau$ in $\mathcal{D}$ with rewards and terminations for each downstream task based on the propositional labels in $\mathcal{D}$, then trained the model to directly predict these quantities. Optimal value estimates were trained via offline RL in a similar manner as GCR. Bespoke Behaviour Cloning (BC) is a neural network policy that directly imitates successful trajectories in $\mathcal{D}$ for every downstream task. Due to the limited number of reward-worthy trajectories in $\mathcal{D}$, we considered any trajectory achieving positive return on that task as successful.
To address RQ4, we compared various reward shaping schemes for GCR. No RS uses predicted RM rewards without any reward shaping (i.e. Algorithm 1). High-Level RS, inspired by Camacho et al. [46], uses a potential function that only considers the current RM state, but not the current MDP state.
Hyperparameters were tuned via a line search over batch size, learning rate, L1 regularization coefficient, and epochs (in that order), using a held-out 10% of the trajectories in $\mathcal{D}$. Final models were retrained on the full data. Full details of the experimental setup can be found in Appendix B.3.
7.3 Results
For each method (excluding behaviour cloning, which directly yields a policy), and for each task in Table 1, we trained a policy from scratch using PPO [73] (training details are reported in Appendix B.3). Recall that the agent does not observe ground-truth rewards or $\mathcal{L}^{*}$, and must generate its own learning signals. We report the performance of each policy at the end of training with respect to $\mathcal{L}^{*}$ (i.e. the ground-truth labelling function used to label $\mathcal{D}$) in Table 2. We measured performance as undiscounted return averaged over 100 episodes for GeoGrid or 20 episodes for DrawerWorld, over five independent RL runs. In Appendix B.4, we report RL learning curves, both with respect to the agent’s own reward model (without shaping rewards) and ground-truth rewards under $\mathcal{L}^{*}$.
GCR consistently solves novel RM tasks. Results show that GCR is the only method that consistently learns successful policies across both GeoGrid and DrawerWorld. It significantly outperforms even the best trajectories in $\mathcal{D}$ on Loop, Lift-Red-Box and Pickup-Each-Box (for the latter task, no successful demonstrations exist in $\mathcal{D}$). We attribute this to GCR’s ability to reuse its learned knowledge towards novel task compositions.
Non-compositional methods fail to generalize. Both LTL-BC, which is trained to imitate trajectories conditioned on entire LTL specifications, and bespoke methods, which directly predict task-specific rewards or behaviours given advance knowledge of the downstream tasks, fail to consistently produce reward-worthy behaviours. The learning curves show that Bespoke Reward Model assigns near-zero rewards to most trajectories in GeoGrid—likely due to the rarity of positive demonstrations in $\mathcal{D}$. In DrawerWorld, it produces highly misaligned rewards, resulting in a policy that achieves high return under its own reward model, but near-zero return under $\mathcal{L}^{*}$.
Reward shaping enables RL in sparse domains. Potential-based shaping provides modest improvements in GeoGrid, but is critical for success in DrawerWorld, where behaviours like opening a drawer and picking up a box are nearly impossible to discover with random exploration. The agent fails to learn under alternative reward shaping schemes (No RS, High-Level RS).
Table 2: Comparison of methods when used to train an RL policy on novel tasks using self-generated rewards. We report performance (undiscounted return, with respect to a ground-truth interpretation of each RM) at the end of training, averaged over 5 seeds with standard error.
Task GCR (Ours) LTL-BC Bespoke Bespoke No RS High-Level RS
Reward Model BC
GeoGrid
Sequence $\textbf{1.00}\pm 0.00$ $0.04\pm 0.01$ $0\pm 0$ $0.05\pm 0.01$ $0.94\pm 0.03$ $\textbf{1.00}\pm 0.00$
Loop $\textbf{5.36}\pm 0.08$ $0.03\pm 0.01$ $0\pm 0$ $0.04\pm 0.01$ $4.68\pm 0.05$ $5.27\pm 0.08$
Logic $\textbf{0.94}\pm 0.01$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $00.00\pm 0$ 0.00 $\textbf{0.94}\pm 0.01$
Safety $\textbf{1.00}\pm 0.00$ $-0.84\pm 0.01$ $-0.14\pm 0.01$ $-0.85\pm 0.01$ $0.23\pm 0.11$ $0.97\pm 0.01$
DrawerWorld
Lift-Red-Box $\textbf{1538}\pm 130$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$
Pickup-Each-Box $\textbf{1.00}\pm 0.00$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$
Show-Green-Box $\textbf{0.61}\pm 0.06$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$
8 Discussion and Limitations
In reflection of RQ1-RQ4, our experiments demonstrate the following advantages of our proposed Ground-Compose-Reinforce (GCR) approach. RQ1: GCR learns interpretations of formal language from data that are better aligned with the ground truth vs. end-to-end grounding approaches. RQ2: GCR generalizes to out-of-distribution tasks and behaviours that are rarely or never observed in the grounding dataset $\mathcal{D}$. RQ3: GCR allows an agent to learn new tasks via RL without supervision. RQ4: GCR with potential-based reward shaping effectively learns RL tasks in sparse settings. We now discuss the main limitations of this work, and opportunities for future work.
Extension to Other Languages. We study the GCR framework from the perspective of Reward Machines, however, we believe our core ideas apply to any formal language that admits compositional structure, such as those that deal with objects [74] and relations [75]. Extending our insights to systems interfaced via natural language (NL), which is often argued to be compositional [76], presents an important challenge. One promising approach is to autoformalize NL into formal language as an intermediate representation [34, 35, 36, 37]. In Appendix Section C, we show how modern LLMs like ChatGPT-4o and o3 (which do not need to be grounded in our specific environments) can zero-shot autoformalize NL reward function descriptions into RMs.
@@ 8 Discussion and Limitations
Extension to Other Problem Settings. Grounded language acquisition is a prerequisite for a myriad of language-conditioned problem settings. While we explore how grounded language can be leveraged for tasking RL agents without external reward supervision, future work could extend our insights to zero-shot execution of language tasks [5, 38, 44], question answering [77, 78], and interactive task learning [79, 80].
@@
Reward Hacking. Misalignment between the agent’s interpretation of language and human intent can lead to unintended or harmful behaviours, posing potential societal risks if deployed. We aim to mitigate this through the use of a formal language, which is unambiguous over its propositional vocabulary. Still, ambiguity in the propositions themselves remains a concern. Future work could explore its impact, though prior works suggest the structure of formal languages may support more transparent reasoning under such uncertainty [55, 56].
Assumptions on Grounding Dataset $\mathcal{D}$. We assume the grounding dataset $\mathcal{D}$ is labelled with values of propositions. Future work could explore practical and scalable labelling methods, such as crowdsourced annotations [81] or self-supervised learning [82]. Our motivation also stems from the growing need for data-efficient methods of building AI systems with general capabilities. We show that even with limited data, our agent generalizes to diverse out-of-distribution tasks, a step toward building systems that require fewer resources, and lower environmental cost.
@@ 9 Conclusion
We present Ground-Compose-Reinforce, a framework for tasking RL agents directly via formal language. We propose a bottom-up approach to grounding language—first learning the meanings of individual symbols from data, then using formal compositional semantics to interpret complex tasks. We implement Ground-Compose-Reinforce for Reward Machines and show how any Reward Machine task can be grounded in dense, shaped rewards from a small grounding dataset of trajectories. Our experiments on challenging, sparse-reward tasks demonstrate its superior ability to interpret and solve novel tasks compared to grounding language directly from end to end.

### v2

@@ 3. RQ3
    Propositional Sparsity: Can the agent operate in extremely long-horizon environments where propositional values are hard to alter with random exploration?
@@ 7.1 Experimental Setup
Figure 4: DrawerWorld is a custom Meta-World environment where the agent can interact with two drawers and three boxes. Propositions capture whether: each drawer is open; each box is lifted by the agent; a given box is in a given drawer.
@@
We considered two domains: an image-based gridworld with randomized object locations called GeoGrid (introduced in the running example), and a Meta-World-based robotics environment called DrawerWorld (Figure 4). Full details on the setup can be found in Appendix B.1.
Pretraining Datasets. We collected $\mathcal{D}$ under minimal assumptions about downstream tasks. In GeoGrid, $\mathcal{D}$ contains 5000 trajectories generated by a random policy. In DrawerWorld, we manually operated the robot to collect 350 trajectories involving generic behaviours (e.g., opening drawers, lifting boxes). To evaluate OOD generalization, we constrained $\mathcal{D}$ to only contain trajectories interacting with at most one box in DrawerWorld. Finally, trajectories were labelled using a handcrafted labelling function.
Tasks. We designed a diverse set of RM tasks (Table 1) that target behaviours rarely or never seen in $\mathcal{D}$. The GeoGrid tasks evaluate whether the agent can produce fine-tuned behaviours beyond the random-action trajectories observed in $\mathcal{D}$. The DrawerWorld tasks evaluate whether the agent can solve complex manipulation tasks that require composing behaviours observed in $\mathcal{D}$ (e.g., Pickup-Each-Box requires handling all three boxes, while trajectories in $\mathcal{D}$ interact with at most one box).
7.2 Method and Baselines
We benchmarked Ground-Compose-Reinforce (GCR) against several non-compositional baselines. Methods based on online RL (GCR, Bespoke Reward Model) use PPO [73] to train a policy from scratch. During execution, GCR captures memory via the RM state while all non-RM-based baselines encode the observation history using an additional GRU [74]. See Appendix B.2 for full implementation details and Appendix B.3 for full training details.
Ground-Compose-Reinforce (ours). We implemented GCR with potential-based reward shaping as described in Sections 5 and 6. Both the predicted labelling function $\hat{\mathcal{L}}$ and PVFs are neural networks trained on $\mathcal{D}$ via supervised learning and offline RL, respectively.
LTL-conditioned Behaviour Cloning (LTL-BC) is a neural network policy $\pi(a_{t}|h_{t},\varphi)$ that directly maps LTL specifications $\varphi$ to behaviours. We labelled each trajectory $\tau$ in $\mathcal{D}$ with an LTL description $\varphi$ (based on the propositional labels for $\tau$), then trained the policy to maximize the log-likelihood of actions in $\tau$, conditioned on the history $h_{t}$ and $\varphi$. For each downstream task in Table 1, we prompted the policy with an LTL formula that aligns with that task.
We also trained bespoke models with advance knowledge of the downstream RM tasks. Bespoke Reward Model directly predicts rewards, optimal values, and terminations for all downstream tasks simultaneously. We labelled each trajectory in $\mathcal{D}$ with ground-truth rewards and terminations for each task based on the propositional labels, then trained the model to directly predict these quantities given the history $h_{t}$. Value estimates were trained via offline RL in a similar manner to GCR. Finally, we trained a policy via RL while using the learned value function for potential-based reward shaping. Bespoke Behaviour Cloning (BC) is a neural network policy that directly imitates successful trajectories in $\mathcal{D}$ for every downstream task. Due to the limited number of reward-worthy trajectories in $\mathcal{D}$, we considered any trajectory achieving positive return on that task as successful.
@@ 7.2 Method and Baselines
Table 1: List of RM tasks. For each, we report the mean ($\mu_{\mathcal{D}}$) and max ($\max_{\mathcal{D}}$) undiscounted return over trajectories in $\mathcal{D}$, along with the max achievable expected return of any policy (Max; if unknown, we report the highest average return observed in our experiments). Some tasks involve behaviours that are rarely or never observed in $\mathcal{D}$.
Task Description Return
GeoGrid $\mu_{\mathcal{D}}$ $\mathrm{max}_{\mathcal{D}}$ Max
Sequence Go to a red $\triangle$, then a green $\triangle$, then a blue $\triangle$. $0.04$ 11 11
Loop Repeatedly go to a red $\triangle$, then a green $\triangle$, then a blue $\triangle$. $0.04$ 33 $5.36$
Logic Go to all six objects, but always go to red objects before blue objects, and blue objects before green objects. $0.00$ 11 11
Safety Go to a red object, then a blue object, then a green object, but always avoid $\triangle$. 0.84\!-\!0.84 11 11
DrawerWorld
Hold-Red-Box Lift and hold the red box as long as possible. $41.7$ $736$ $1538$
Pickup-Each-Box Pick up the red box, then the blue box, then the green box. 0 0 11
Show-Green-Box Reveal the green box if it’s in a closed drawer, then lift it. $0.22$ 11 11
We also compared various reward shaping schemes for GCR. No RS directly uses RM rewards without any reward shaping (i.e. Algorithm 1). High-Level RS, inspired by Camacho et al. [34], uses a potential function that only considers the current RM state, but not the current MDP state.
7.3 Results
We ran each method’s training pipeline five times and report the average final performances in Table 2. Performance was measured by undiscounted return (averaged over 100 evaluation episodes for GeoGrid or 20 for DrawerWorld), where rewards are with respect to the ground-truth labelling function $\mathcal{L}^{*}$. In Appendix B.4, we report RL learning curves, both with respect to the agent’s own reward model (without shaping rewards) and ground-truth rewards under $\mathcal{L}^{*}$.
GCR consistently solves novel tasks, even when no successful demonstrations exist in $\mathcal{D}$. On Loop, Hold-Red-Box and Pickup-Each-Box, it significantly outperforms even the best trajectories in $\mathcal{D}$. We attribute this success to GCR’s ability to learn transferable knowledge from $\mathcal{D}$ and apply it towards novel task compositions, while fine-tuning behaviours with (self-supervised) RL.
All non-compositional baselines fail to reliably solve any task. Results show that LTL-BC, Bespoke Reward Model, and Bespoke BC do not fare well with limited pretraining trajectories. Learning curves show that the Bespoke Reward Model assigns near-zero rewards to most trajectories in GeoGrid, likely due to the rarity of positive demonstrations in $\mathcal{D}$. In DrawerWorld, it produces misaligned rewards, leading to reward hacking (evidenced by high returns under the learned reward model but low returns under the ground-truth $\mathcal{L}^{*}$).
Reward shaping enables long-horizon RL. Our reward shaping strategy yields modest improvements in GeoGrid, but is critical to success in DrawerWorld, where behaviours like opening a drawer and picking up a box are nearly impossible to discover from random exploration alone.
We conclude the following. GCR faithfully elicits behaviours from RM specifications, outperforming non-compositional approaches (RQ1). Moreover, GCR compositionally generalizes to OOD behaviours beyond those observed in $\mathcal{D}$ (RQ2). Finally, our compositional reward shaping strategy for GCR enables RL in long-horizon settings involving propositional sparsity (RQ3).
Table 2: Comparison of methods for eliciting behaviours from high-level task specifications. We report performance (undiscounted return with respect to ground-truth rewards) averaged over 5 runs with standard error.
Task GCR (Ours) LTL-BC Bespoke Bespoke GCR (Ours) GCR (Ours)
Reward Model BC No RS High-Level RS
@@ 7.3 Results
Sequence $\textbf{1.00}\pm 0.00$ $4\pm 0.01 0±00\pm 0$ $0.05\pm 0.01$ $0.94\pm 0.03$ $\textbf{1.00}\pm 0.00$
Loop $\textbf{5.36}\pm 0.08$ $3\pm 0.01 0±00\pm 0$ $0.04\pm 0.01$ $4.68\pm 0.05$ $5.27\pm 0.08$
Logic $\textbf{0.94}\pm 0.01$ $\pm 0 0±0$$\pm 0$ $\pm 0 0.0$$00\pm 0.00$ $\textbf{0.94}\pm 0.01$
Safety $\textbf{1.00}\pm 0.00$ $0.84\pm 0.01$ $0.14\pm 0.01$ $0.85\pm 0.01$ $0.23\pm 0.11$ $0.97\pm 0.01$
@@
Hold-Red-Box $\textbf{1538}\pm 130$ $\pm 0 0±0$$\pm 0$ $\pm 0 0±0$$\pm 0$ $0\pm 0$
Pickup-Each-Box $\textbf{1.00}\pm 0.00$ $\pm 0 0±0$$\pm 0$ $\pm 0 0±0$$\pm 0$ $0\pm 0$
Show-Green-Box $\textbf{0.61}\pm 0.06$ $\pm 0 0±0$$\pm 0$ $\pm 0 0±0$$\pm 0$ $0\pm 0$
7.4 Extending Ground-Compose-Reinforce with a Natural Language Interface
While natural language (NL) is often argued to have compositional properties [75], exploiting this compositionality in agentic language models (e.g. vision-language-action models) remains an open challenge. In Appendix C, we show that our GeoGrid RMs can be autoformalized directly from an NL reward function description using OpenAI’s o3 model, zero-shot—i.e., without fine-tuning on trajectories or other forms of grounding in our specific environments. Thus, we posit that leveraging compositional representations like RMs can be an effective way of building NL-interfaced agents in settings with limited labelled trajectory data (i.e. where $|\mathcal{D}|$ is small).
8 Future Work and Limitations
Extension to Other Compositional Representations: In this work, we propose an end-to-end framework for grounding high-level specifications in behaviours that leverages the compositionality inherent in RMs. However, we believe our core insights apply to a wide range of compositional representations such as those that deal with objects [76] and relations [77].
@@ 8 Future Work and Limitations
Extension to Other Problem Settings: Grounding language is a prerequisite for a myriad of language-conditioned problem settings. We consider an “RL-in-the-loop” setting, but future works could extend our insights to zero-shot execution of language tasks [5, 40, 46], question answering [78, 79], and interactive task learning [80, 81].
@@
Reward Hacking: Misalignment between an agent’s interpretation of a task and human intent can lead to harmful consequences, particularly in RL [82]. The use of formal specifications like RMs, which are unambiguous over the propositional vocabulary, can partially mitigate this, but ambiguity in the propositions themselves remains a concern in Ground-Compose-Reinforce. Prior works suggest that RM structure can be exploited to improve decision making under such ambiguity [55, 56].
Assumptions on $\mathcal{D}$: We assume that trajectories in $\mathcal{D}$ are labelled with values for a fixed set of propositions. Future works could explore other representations of propositions (e.g. as text) as well as scalable labelling methods (e.g. crowdsourced annotations [69] or self-supervised learning [70]).
@@ 9 Conclusion
This work presents Ground-Compose-Reinforce, an end-to-end framework for training RL agents directly from Reward Machine specifications—without oracle reward or labelling functions. A key challenge that we address is grounding these high-level task specifications in executable behaviours, given an agent’s perception and action capabilities. We find that exploiting compositional task structure is critical to faithfully capturing this grounding from limited data. Starting from only 350 labelled pretraining trajectories, we show that our technical approach scales to temporally extended manipulation tasks in Meta-World while generalizing out-of-distribution to behaviours that never appear in pretraining. Moreover, we show that in some cases, Reward Machines can be autoformalized directly from natural language reward function descriptions to expose this temporal task structure.
More broadly, we show that leveraging language compositionality presents a promising pathway to building language-driven agents without relying on massive language-labelled data. Future work could explore the extension of these ideas to large-scale agentic language models such as vision-language-action models.